# Minimal Thermodynamic Cost of Computing with Circuits

### v2

@@ III.1 Dynamics of the circuit
We assume input values are jointly sampled from an input distribution $p_{in}(\boldsymbol{x}_{in})$. The dependency of a gate on its parents is expressed by the conditional distribution $\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})$ for each non-input gate $\mu\in V_{nin}$. The conditional distributions $\pi_{\mu}$ together specify the conditional probability of the state of the non-input gates given the states of input gates:
$\mathrm{R}(\boldsymbol{x}_{nin}|\boldsymbol{x}_{in})=\prod_{\mu\in V/V_{in}}\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})$ (5)
Note that, regardless of the gate being deterministic or noisy, since the inputs are randomly sampled from an input distribution $p_{in}$, the state of each gate, and hence the joint state of the circuit is a random variable. After the end of a complete run and before the beginning of the next run, the joint state of the entire circuit has a following probability distribution (see Fig. 3A)
$p^{0}(\boldsymbol{x})=R(\boldsymbol{x}_{nin}|\boldsymbol{x}_{in})p_{in}(\boldsymbol{x}_{in})$ (6)
@@
$p_{\mu:}(\boldsymbol{x}_{:\mu})$ Marginal distribution of subset of gates following gate $\mu$
@@
$\mathcal{MC}_{\mu}$ MMC of running gate $\mu$ in a circuit
$\mathcal{MC}_{l}$ MMC of running layer $l$ in a circuit
$\mathcal{MC}_{ow}(p_{in})$ MMC of overwriting inputs with input distribution $p_{in}$
$\mathcal{MC}(C_{n},p_{in})$ Total MMC of running the circuit $C_{n}$ for the input distribution $p_{in}$
@@ III.1.1 Notation
Analogously, $p_{in}$ and $p_{nin}$ are used to denote the joint distribution over the states of input nodes and non-input gates respectively.
$p_{nin}(\boldsymbol{x}_{nin})=\sum_{x_{in}}p^{0}(\boldsymbol{x})$ (9)
@@ III.1.2 Joint distribution after the re-initialization of the input nodes
$p^{1}(\boldsymbol{x})=p_{in}(\boldsymbol{x}_{in})p_{nin}(\boldsymbol{x}_{nin}).$ (10)
@@ III.1.3 Joint distribution after the update of gate $\mu$
As mentioned earlier and depicted in Fig. 2, before gate $\mu$ updates, $\boldsymbol{x}_{\mu}$ is correlated with $\boldsymbol{x}_{\mu:}$ but independent of $\boldsymbol{x}_{:\mu}$. Therefore, the joint distribution before the update of gate $\mu$ is
@@
$p(\boldsymbol{x}_{:\mu})p(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:})\to p(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p(\boldsymbol{x}_{\mu:})$ (13)
III.2 MMC of re-initializing inputs with new values.
Figure 3: Evolution of joint distribution under re-initialization of inputs and re-running the circuit gate-by-gate. (A) After a complete run, the joint distribution reflects maximal correlation among all gates. This correlation arises from the initial input distribution and the network of dependencies within the circuit. (B) In the next run, the input nodes are re-sampled from the input distribution $p(\boldsymbol{x}_{in})$, making the new inputs independent of the states of non-input gates from the previous run. As a result, the joint distribution after re-initialization is factored as $p(\boldsymbol{x}_{in})p(x_{1},...,x_{5})$, indicating that inputs and non-input gates are now statistically independent. (C) When gate $x_{1}$ updates based on the newly re-sampled input values, it becomes correlated with the input nodes while simultaneously losing correlation with the rest of the gates. Consequently, the joint distribution evolves to $p(\boldsymbol{x}_{in},x_{1})p(x_{2},...,x_{5})$. Similarly, (D), (E), (F), and (G) illustrate the sequential updates of the remaining gates and the corresponding evolution of the distribution. In particular, (G) demonstrates that after a complete run of the circuit, the distribution returns to its initial form as shown in (A).
During the process of overwriting new input values, the input nodes evolve independently of the rest of the non-input nodes which remain unchanged. Therefore, updating input nodes with new values is a sub-system process where the sub-system $\boldsymbol{x}_{in}$ evolves independently of $\boldsymbol{x}_{nin}$ while the later remains unchanged. The prior distribution for this subsystem process is product distribution, expressed as $q_{0}(\boldsymbol{x}_{in})q_{0}(\boldsymbol{x}_{nin})$. Since new values of the input nodes are sampled from $p_{in}(x_{in})$, this prior distribution evolves to $p_{in}(\boldsymbol{x}_{in})q_{0}(\boldsymbol{x}_{nin})$.
$q_{in}(\boldsymbol{x}_{in})q_{nin}(\boldsymbol{x}_{nin})\longrightarrow p_{in}(\boldsymbol{x}_{in})q_{nin}(\boldsymbol{x}_{nin})$ (14)
while the actual distribution evolves from $p^{0}(\boldsymbol{x})$ to $p_{in}(\boldsymbol{x}_{in})p_{nin}(\boldsymbol{x}_{nin})$:
$p^{0}(\boldsymbol{x})\longrightarrow p_{in}(\boldsymbol{x}_{in})p_{nin}(\boldsymbol{x}_{nin})$ (15)
Since new input values are totally independent of the states of the rest of the gates, $\mathrm{I}_{1}(\boldsymbol{x}_{in};\boldsymbol{x}_{nin})=0$, and the drop in mutual information is
$\Delta\mathrm{I}(\mathrm{X}_{in};\mathrm{X}_{nin})$ $=\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{nin})-\mathrm{I}_{1}(\mathrm{X}_{in};\mathrm{X}_{nin})$ (16)
$=\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{nin})$ (17)
The MMC of overwriting the input values is
@@
$=$ $\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{nin})+D(p_{in}||q_{in})$ (19)
Moreover, if $\mathrm{X}_{1}$ is a deterministic function of $\mathrm{X}_{in}$, then $\mathrm{I}_{0}(\mathrm{X}_{in};\mathrm{X}_{1})=S(p_{in})$. In that case, the overwriting MMC is given by
$\mathcal{MC}_{ow}=\mathcal{C}~{}(p_{in},q_{in})$ (20)
where $\mathcal{C}~{}(p_{in},q_{in})$ of $q_{in}$ relative to $p_{in}$.
III.3 MMC of gate-by-gate or layer-by-layer implementation of the circuit.
@@
Figure 4: Layer-by-layer implementation of a circuit. (A) State of the circuit after a complete run. (B) Updating of input nodes for the next run, with values re-sampled from $p(\boldsymbol{x}_{in})$. (C) The first layer consists of gates $x_{1}$ and $x_{2}$. As they update together based on the newly sampled input values, they become correlated with the input nodes while losing correlation with the rest of the gates. Consequently, the joint distribution evolves to $p(\boldsymbol{x}_{in},x_{1},x_{2})p(x_{3},x_{4},x_{5})$. (D) Gates $x_{3}$ and $x_{4}$ make up layer 2. As they update based on the new values of gates $x_{1}$ and $x_{2}$, they become correlated with them and the input nodes, while becoming independent of the state of $x_{5}$. The distribution then evolves to $p(\boldsymbol{x}_{in},x_{1},...,x_{4})p(x_{5})$. Finally, in (E), gate $x_{5}$, constituting layer 3, is updated, and the distribution returns to its initial form as shown in (A).
The MMC of updating the gate $\mu$ is given by,
@@
Using equations (11), (12), (21), and (22), the expression for the MMC $\mathcal{MC}_{\mu}$ simplifies to
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)})\right)-D\left(p_{\mathrm{pa}(\mu)})||q_{\mathrm{pa}(\mu)}\right),$ (24)
where $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$ denotes the mutual information between gate $\mu$ and its children $\mathrm{ch}(\mu)$ prior to the update of $\mu$. Specifically, the mutual information is evaluated with respect to the distribution $p^{\mu}(\boldsymbol{x})$, which characterizes the state of the system immediately before the update of $\mu$. The derivation of Eq. (24) is provided in App. A.3. The total MMC of executing the entire circuit through a sequential gate-by-gate implementation is given by the sum of the MMCs of individual gates and the MMCs incurred when updating input nodes,
$\mathcal{MC}(C_{n},p_{in})=$ $\;\mathcal{MC}_{ow}+\sum_{\mu\in V\backslash V_{in}}\mathcal{MC}_{\mu}$ (25)
$=$ $\;\mathcal{C}~{}(p_{in},q_{in})+\sum_{\mu\in V\backslash V_{in}}\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+\Delta_{\mu}.$ (26)
where
$\Delta_{\mu}:=D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)})\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right).$ (27)
@@
It is straightforward to extend the analysis for a layer-by-layer implementation of circuit (see Fig. 4). Based on a topological ordering of a DAG, it is possible to stratify the set of gates in a natural way into layers. Starting with the set of input nodes $V_{0}:=\{\mu\in V:\mathrm{pa}(\mu)=\emptyset\}$, the set of gates constituting the $l$-th layer is defined as,
$V_{l}:=\{\mu\in V\backslash(V_{0}\cup...\cup V_{l-1}):\mathrm{pa}(\mu)\cap(V_{1}\cup...\cup V_{l-1})\neq\emptyset\},$ (28)
@@
for any $l\in\{1,...,L\}$. Let $X_{l}$ be the random variable denoting the joint state of gates in layer $l$ and let $\boldsymbol{x}_{l}$ denote a value of $X_{l}$. In a layer-by-layer implementation, all the gates in a layer are updated simultaneously. Then, analogous to Eq. 24, the MMC of updating layer $l$ in the circuit is given by,
$\mathcal{MC}_{l}=$ $\mathrm{I}_{l}(\mathrm{X}_{l};\mathrm{X}_{\mathrm{ch}(l)})$ (29)
$+D\left(p_{l}p_{\mathrm{pa}(l)}||q_{l,\mathrm{pa}(l)}\right)-D\left(p_{\mathrm{pa}(l)}||q_{\mathrm{pa}(l)}\right),$
@@
where $p_{l}$, $p_{\mathrm{pa}(l)}$, $q_{l,\mathrm{pa}(l)}$, and $q_{\mathrm{pa}(l)}$ are defined for layer $l$ and have their usual meaning as in Eq. 24. The derivation of Eq. 29 is provided in App. A.3.
@@
Updating multiple gates simultaneously in a layer-by-layer approach, instead of updating gates sequentially one by one, represents a form of time-coarse graining. This is because grouping several updates into a single time step effectively reduces the temporal resolution of the system’s dynamics. As demonstrated in yadav2024mismatch , MMC decreases under time-coarse graining. Applied to circuits, this result implies that the MMC of a layer-by-layer implementation always serves as a lower bound on the MMC of a gate-by-gate implementation. In Fig. 5, we compare the MMCs of the layer-by-layer and gate-by-gate implementations of the same ripple-carry adder. The gate-by-gate MMC is computed using Eq. 24, while the layer-by-layer MMC is computed using Eq. 29. As a consequence of the lowering in MMC under time-coarse graining, the layer-by-layer MMC is lower than the gate-by-gate MMC for the same circuit, across ripple-carry adder circuits with varying input lengths.
@@
Additionally, note that as the computation in a circuit progress—whether gate-by-gate or layer-by-layer—the distribution over the states cycles through a sequence of distributions (see Fig. 3 and 4). It is important to emphasize that, for the calculation of the total MMC, when calculating the total MMC, the phase at which the sequence of distributions is started does not impact the final MMC, as long as the full cycle of execution is completed.
@@
Moreover, the sequence of gate implementations in a circuit can be modeled as a periodic process. In practical computational devices based on circuits, the execution of gates or layers is typically driven by a clock. From this perspective, the process is periodic, with gates executed at each clock cycle, sequentially leading to the completion of the full circuit. In Appendix B.3, we present the MMC calculation for this clock-driven model of circuit execution. This approach yields the same expression for the MMC as given in Eq. 23.
@@
In the next section, we use Eq 24 to calculate the MMC of various circuit families and broadly investigate how does the circuit structure affect their MMC. In particular, we define MMC complexity, analogous to the size and depth complexity of a circuit family. We will then establish results that relate MMC complexity to size complexity.
@@
Figure 5: Comparison of layer-by-layer and gate-by-gate mismatch costs for the ripple-carry adder circuit family. As the circuit size increases with the input size, the MMC grows for both the layer-by-layer (dashed red) and gate-by-gate (solid blue) implementations. Notably, the layer-by-layer MMC remains consistently lower than the gate-by-gate MMC. In this comparison, the input distribution is uniformly random for both implementations, and the priors of all gates in the circuits are also uniform.

### v3

@@ III.1 Dynamics of the circuit
We assume input values are jointly sampled from an input distribution $p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})$. We assume each run of the circuit is evaluated on the entire input distribution $p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})$. The dependency of a gate on its parents is expressed by the conditional distribution $\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})$ for each non-input gate $\mu\in V_{\mathrm{nin}}$. The conditional distributions $\pi_{\mu}$ together specify the conditional probability of the state of the non-input gates given the states of input gates:
$\mathrm{R}(\boldsymbol{x}_{\mathrm{nin}}|\boldsymbol{x}_{\mathrm{in}})=\prod_{\mu\in V/V_{\mathrm{in}}}\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})$ (5)
Note that, regardless of the gate being deterministic or noisy, since the inputs are randomly sampled from an input distribution $p_{\mathrm{in}}$, the state of each gate, and hence the joint state of the circuit is a random variable. After the end of a complete run and before the beginning of the next run, the joint state of the entire circuit has a following probability distribution (see Fig. 3A)
$p^{0}(\boldsymbol{x})=R(\boldsymbol{x}_{\mathrm{nin}}|\boldsymbol{x}_{\mathrm{in}})p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})$ (6)
@@
$\boldsymbol{x}_{\mu:}$ Joint state of gates following gate $\mu$
$\boldsymbol{x}_{in}$ Joint state of the inputs
$\boldsymbol{x}_{nin}$ Joint state of non inputs gates
$\boldsymbol{x}_{-\mu}$ Joint state of all gates except gate $\mu$ and the parents of gate $\mu$
@@
$p_{\mu:}(\boldsymbol{x}_{\mu:})$ Marginal distribution of subset of gates following gate $\mu$
@@
$\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})$ Mutual information between gate $\mu$ and its parents $\mathrm{ch}(\mu)$ with reference to distribution $p^{\mu}(\boldsymbol{x})$
$\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mu:})$ Mutual information between gate $\mu$ and all gates following $\mu$ with reference to distribution $p^{\mu}(\boldsymbol{x})$
$\mathcal{MC}_{\mu}$ Mismatch cost of running gate $\mu$ in a circuit
$\mathcal{MC}_{l}$ Mismatch cost of running layer $l$ in a circuit
$\mathcal{MC}_{ow}(p_{\mathrm{in}})$ Mismatch cost of overwriting inputs with input distribution $p_{\mathrm{in}}$
$\mathcal{MC}(C_{n},p_{\mathrm{in}})$ Total mismatch cost of running the circuit $C_{n}$ for the input distribution $p_{\mathrm{in}}$
@@ III.1.1 Notation
Analogously, $p_{\mathrm{in}}$ and $p_{\mathrm{nin}}$ are used to denote the joint distribution over the states of input nodes and non-input gates respectively.
$p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})=\sum_{x_{\mathrm{in}}}p^{0}(\boldsymbol{x}),\quad p_{\mathrm{nin}}(\boldsymbol{x}_{\mathrm{nin}})=\sum_{x_{\mathrm{in}}}p^{0}(\boldsymbol{x}).$ (9)
A more comprehensive list of notation is provided in Table 1.
@@ III.1.2 Joint distribution after the re-initialization of the input nodes
$p^{1}(\boldsymbol{x})=p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})p_{\mathrm{nin}}(\boldsymbol{x}_{\mathrm{nin}}).$ (10)
@@ III.1.3 Joint distribution after the update of gate $\mu$
Before gate $\mu$ updates, its state $\boldsymbol{x}_{\mu}$ is correlated with $\boldsymbol{x}_{\mu:}$ from the previous run of the circuit and it is uncorrelated with $\boldsymbol{x}_{:\mu}$ because of the new values. The joint distribution before the gate $\mu$ updates has the following form:
@@
This transformation of probability distribution upon executing gate $\mu$ can be represented by
$p(\boldsymbol{x}_{:\mu})p(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:})\to p(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p(\boldsymbol{x}_{\mu:}).$ (13)
Running gate $\mu$ correlates its output with its parents while making it conditionally independent of downstream variables. Intuitively, executing gate $\mu$ redistributes correlations: it introduces correlations between the gate’s output and its parents, while breaking correlations with downstream variables. This type of redistribution of correlations occurs at each non-input gate as the circuit is evaluated in topological order, continually creating and erasing statistical dependencies among subsets of variables.
III.2 Mismatch cost of re-initializing inputs with new values.
During the process of overwriting new input values, the input nodes evolve independently of the rest of the non-input nodes which remain unchanged. Therefore, updating input nodes with new values is a sub-system process where the sub-system $\boldsymbol{x}_{\mathrm{in}}$ evolves independently of $\boldsymbol{x}_{\mathrm{nin}}$ while the latter remains unchanged. For such a sub-system process, the prior distribution is product distribution, expressed as $q_{0}(\boldsymbol{x}_{\mathrm{in}},\boldsymbol{x}_{\mathrm{nin}})=q_{0}(\boldsymbol{x}_{\mathrm{in}})q_{0}(\boldsymbol{x}_{\mathrm{nin}})$ (see App. A.1). Since new values of the input nodes are sampled from $p_{\mathrm{in}}(x_{\mathrm{in}})$, this prior distribution evolves to $\tilde{q}_{0}(\boldsymbol{x}_{\mathrm{in}},\boldsymbol{x}_{\mathrm{nin}})p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})q_{0}(\boldsymbol{x}_{\mathrm{nin}})$.
$q_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})q_{\mathrm{nin}}(\boldsymbol{x}_{\mathrm{nin}})\longrightarrow p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})q_{\mathrm{nin}}(\boldsymbol{x}_{\mathrm{nin}})$ (14)
while the actual distribution evolves from $p^{0}(\boldsymbol{x})$ to $p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})p_{\mathrm{nin}}(\boldsymbol{x}_{\mathrm{nin}})$:
$p^{0}(\boldsymbol{x}_{\mathrm{in}},\boldsymbol{x}_{\mathrm{nin}})\longrightarrow p_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})p_{\mathrm{nin}}(\boldsymbol{x}_{\mathrm{nin}})$ (15)
Since new input values are totally independent of the states of the rest of the gates, $\mathrm{I}_{1}(\boldsymbol{x}_{\mathrm{in}};\boldsymbol{x}_{\mathrm{nin}})=0$, and the drop in mutual information is
$\Delta\mathrm{I}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})$ $=\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})-\mathrm{I}_{1}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})$ (16)
$=\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})$ (17)
Using Eq. (14) and (15), the mismatch cost of overwriting the input values evaluates to
@@
$=$ $\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})+D(p_{\mathrm{in}}||q_{\mathrm{in}}),$ (19)
where the mutual information term $\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})$ arises because the circuit begins each run in a correlated state: the values stored in the non-input nodes reflect remnants of the previous computation and are therefore statistically dependent on the previous inputs. When the input nodes are re-initialized with fresh values drawn from the new input distribution, these correlations are erased. As a result, the joint distribution over input and non-input nodes becomes a product distribution at the start of the new run, and the loss of these pre-existing correlations results in a term $\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})$ from the change in KL-divergence.
III.3 Mismatch cost of gate-by-gate or layer-by-layer implementation of the circuit.
@@
Figure 4: Layer-by-layer implementation of a circuit. (A) State of the circuit after a complete run. (B) Updating of input nodes for the next run, with values re-sampled from $p(\boldsymbol{x}_{\mathrm{in}})$. (C) The first layer consists of gates $x_{1}$ and $x_{2}$. As they update together based on the newly sampled input values, they become correlated with the input nodes while losing correlation with the rest of the gates. Consequently, the joint distribution evolves to $p(\boldsymbol{x}_{\mathrm{in}},x_{1},x_{2})p(x_{3},x_{4},x_{5})$. (D) Gates $x_{3}$ and $x_{4}$ make up layer 2. As they update based on the new values of gates $x_{1}$ and $x_{2}$, they become correlated with them and the input nodes, while becoming independent of the state of $x_{5}$. The distribution then evolves to $p(\boldsymbol{x}_{\mathrm{in}},x_{1},...,x_{4})p(x_{5})$. Finally, in (E), gate $x_{5}$, constituting layer 3, is updated, and the distribution returns to its initial form as shown in (A).
The mismatch cost of updating the gate $\mu$ is given by,
@@
Using expressions  (11) and  (12) for the actual probability distributions, and  (20) and (21) for the prior distributions, the expressions for the mismatch cost $\mathcal{MC}_{\mu}$ becomes
$\mathcal{MC}_{\mu}=S_{\mu+1}(\mathrm{X})-S_{\mu}(\mathrm{X})+\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right),$ (23)
where $\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})$ denotes the mutual information between $\mu$ and $\mathrm{pa}(\mu)$ after the gate $\mu$ is updated. $S_{\mu}(X)$ and $S_{\mu+1}(X)$ are the entropy of the joint system (the entire circuit) before and after updating gate $\mu$. The derivation of Eq. (23) is provided in App. A.3. More importantly, the total mismatch cost of the entire circuit, given by the sum of mismatch cost of overwriting of inputs (Eq. 19) and mismatch cost running all gates sequentially, turns out to have the following form:
$\mathcal{MC}(C_{n},p_{\mathrm{in}})$ $=\mathcal{MC}_{ow}(p_{\mathrm{in}})+\sum_{\mu\in V_{\mathrm{nin}}}\mathcal{MC}_{\mu}$ (24)
$=D(p_{\mathrm{in}}||q_{\mathrm{in}})+\sum_{\mu\in V_{\mathrm{nin}}}\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\sum_{\mu\in V_{\mathrm{nin}}}\left[D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right)\right].$ (25)
A few remarks about Eq. (25) are in order. In particular, the expression for the MMC of a single gate $\mu$ in Eq. (23) involves the entropy of the full circuit state, which is a global quantity. However, when these contributions are summed over all non-input gates, the intermediate entropy terms telescope and cancel. What remains are only the boundary terms,
$S_{1}(\mathrm{X})-S_{|V_{\mathrm{nin}}|+1}(\mathrm{X}),$ (26)
corresponding to the entropy of the joint circuit state immediately after the inputs are re-initialized and before any non-input gates are executed, and the entropy after all non-input gates have completed their updates. This boundary contribution can be written as
$S_{1}(\mathrm{X})-S_{|V_{\mathrm{nin}}|+1}(\mathrm{X})=-\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}}),$ (27)
reflecting the correlations established between inputs and non-input nodes during the computation. Consequently, after adding the overwriting MMC in Eq. (19), the total mismatch cost simplifies to the expression in Eq. (25).
@@
It is straightforward to extend the analysis for a layer-by-layer implementation of circuit (see Fig. 4). Based on a topological ordering of a DAG, it is possible to stratify the set of gates in a natural way into layers. Starting from the set of input nodes
$V_{0}:={\mu\in V:\mathrm{pa}(\mu)=\emptyset},$ (28)
we define the $l$-th layer recursively by
$V_{l}:=\left\{\mu\in V\setminus\bigcup_{k=0}^{l-1}V_{k}:\mathrm{pa}(\mu)\subseteq\bigcup_{k=0}^{l-1}V_{k}\right\}.$ (29)
This construction assigns each gate to the earliest layer in which all of its parents have already been evaluated, thereby inducing a valid feedforward layering of the circuit.
@@
Let $\mathrm{X}_{l}$ be the random variable denoting the joint state of gates in layer $l$ and let $\boldsymbol{x}_{l}$ denote a value of $\mathrm{X}_{l}$. In a layer-by-layer implementation, all the gates in a layer are updated simultaneously. Let $d$ be the number of layers in the circuit. Then, analogous to Eq. (25), the total mismatch cost of running all non-input layers is given by,
$\mathcal{MC}_{ow}$ $+\sum_{l=1}^{d}\mathcal{MC}_{l}=D(p_{\mathrm{in}}||q_{\mathrm{in}})+\sum_{l=1}^{d}\mathrm{I}_{l+1}(\mathrm{X}_{l};\mathrm{X}_{\mathrm{pa}(l)})$ (30)
$+\sum_{l=1}^{d}\left[D\left(p_{l}p_{\mathrm{pa}(l)}||q_{l,\mathrm{pa}(l)}\right)-D\left(p_{\mathrm{pa}(l)}||q_{\mathrm{pa}(l)}\right)\right],$
@@
where $_{l}, pp$$_{\mathrm{pa}(l)}$, $q_{l,\mathrm{pa}(l)}$, and $q_{\mathrm{pa}(l)}$ are defined for layer $l$ and have their usual meaning as in Eq. (25).
@@
Updating multiple gates simultaneously in a layer-by-layer approach, instead of updating gates sequentially one by one, represents a form of time-coarse graining. This is because grouping several updates into a single time step effectively reduces the temporal resolution of the system’s dynamics. As demonstrated in [17], mismatch cost decreases under time-coarse graining. Applied to circuits, this result implies that the mismatch cost of a layer-by-layer implementation always serves as a lower bound on the mismatch cost of a gate-by-gate implementation. In Fig. 5, we compare the mismatch costs of layer-by-layer and gate-by-gate implementations of a ripple-carry adder. Gate-by-gate mismatch cost is computed using Eq. (25), while the layer-by-layer mismatch cost is computed using Eq. (30). As a consequence of the lowering in mismatch cost under time-coarse graining, the layer-by-layer mismatch cost is lower than the gate-by-gate mismatch cost across ripple-carry adder circuits with varying input lengths.
@@
Additionally, note that as the computation in a circuit progress—whether gate-by-gate or layer-by-layer—the distribution over the states cycles through a sequence of distributions (see Fig. 3 and 4). It is important to emphasize that, for the calculation of the total mismatch cost, when calculating the total mismatch cost, the phase at which the sequence of distributions is started does not impact the final mismatch cost, as long as the full cycle of execution is completed.
@@
Moreover, the sequence of gate implementations in a circuit can be modeled as a periodic process. In practical computational devices based on circuits, the execution of gates or layers is typically driven by a clock [19]. From this perspective, the process is periodic, with gates executed at each clock cycle, sequentially leading to the completion of the full circuit. In Appendix B.3, we present the MMC calculation for this clock-driven model of circuit execution. This approach yields the same expression for the MMC as given in Eq. (22).
@@
In the next section, we use Eq (25) to calculate the mismatch cost of various circuit families and broadly investigate how does the circuit structure affect their mismatch cost. In particular, we define mismatch cost complexity, analogous to the size and depth complexity of a circuit family. We will then establish results that relate mismatch cost complexity to size complexity.
@@
Figure 5: Comparison of layer-by-layer and gate-by-gate mismatch costs for the ripple-carry adder circuit family. Notably, the layer-by-layer mismatch cost remains consistently lower than the gate-by-gate mismatch cost. In this comparison, the input distribution is uniformly random for both implementations, and the priors of all gates in the circuits are also uniform.